OpenStudy (anonymous):

Please help, medal will be given! Evaluate the function rule to find the range of each function for the domain {-3, 0, 5}. y = 3x² − 8 A. {-35, -8, 22} B. {-35, 8, 67} C. {-8, 19, 67} D. {8, 19, 22}

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

c

OpenStudy (michele_laino):

hint: if I replace \(x=-3\) into your formula, I get: \[y = 3 \cdot {3^2} - 8 = 27 - 8 = 19\] so the point \(y=19\) belongs to the range of your function. Now please do the same with \(x=0\) and \(x=5\)